OpenStudy (anonymous):

In "Worked Example" from session 69, INT(cos2x)^2 is given as (x/2) + (sin2x/4) + c. Isn't this incorrect?

OpenStudy (stacey):

I am not looking at that session, but what you have typed is incorrect. The answer you have is for \[\int\limits_{}^{}\cos ^{2}x dx\] \[\int\limits_{}^{}\cos ^{2}2xdx=x/2+(\sin4x)/8+c\]

OpenStudy (anonymous):

yeah I know its incorrect... thats why I asked that... while matching my answer with the one in solution I found this. I thought that the solution must be incorrect. Just needed a double confirmation.